OpenStudy (anonymous):

If the common difference of the A.P. a1, a2, a3 ,.... , is d, then the common difference of the A.P. 2a1+3, 2a2+3, 2a3+3...is ..?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

I would subtract the 1st 2 terms of the second A.P. what do you get ?

OpenStudy (anonymous):

2d+3?

OpenStudy (phi):

do second term minus first term

OpenStudy (anonymous):

2a2-2a1

OpenStudy (phi):

(2a2+3) - ( 2a1+3) = ?

OpenStudy (phi):

two more things: 1) they tell you that a2 - a1 = d (the first A.P.) 2) you can factor 2 out of (2a2 - 2a1)

OpenStudy (phi):

so the difference is 2d

OpenStudy (anonymous):

2d

OpenStudy (anonymous):

ok, i see